# Elias’ Encoding from Lagrangians and Renormalization

### v3

@@ Abstract
In the present paper we give a principled derivation of Elias’ Omega code by combining a constrained variational formulation of prefix coding with a renormalization flow on codeword distributions. Starting from a Lagrangian that minimizes average code length under the Kraft–McMillan constraint, we show that the implied distribution is a fixed point of a coarse–graining map, yielding the canonical iterated logarithm length, up to an additive constant.
This establishes completeness and asymptotic optimality, and connects universal integer coding with coarse–grained entropy, uncertainty–type bounds, and multiplicity relations familiar from statistical physics. The renormalization operator induces a discrete flow that converges to the Elias fixed point for any admissible initialization, up to a bounded error, offering a clean bridge between information–theoretic constraints and RG–style scale invariance.
@@ 1 Introduction
Among several possible encoding procedures for integers, in his paper [1] Elias introduces the so–called $\omega$ encoding. This is a universal prefix code obtained by the simple iterative procedure below. Let $s=\text{bin}(x)$ be the binary string representation of a decimal number $x$, and let $x=\text{dec}(s)$ be the decimal value corresponding to a binary string $s$.
Algorithm 1 Elias’ Omega: Encoding
@@
Algorithm 2 Elias’ Omega: Decoding
@@
Since Shannon’s foundational work [2] and the Kraft–McMillan inequality [3, 4], a central question has been how to construct codes that approach entropy without knowledge of the true source distribution: such codes are called universal. Elias’s $\gamma$, $\delta$, and $\omega$ codes are all universal [1].
In this note, we show that $\omega$ follows directly from a constrained Lagrangian optimization combined with a scale‐invariance requirement for the implied distribution, yielding a clean derivation from fundamental physical and code-theoretic principles.
@@ 2.1 Lagrangian Optimization for Code Length
i.e. a weighted sum of delta-functions together with a continuous probability density. Here $W$ is at most countable, $\rho\geq 0$ is integrable, and
@@
Let us formally recast the Kraft–McMillan inequality [3, 4] for prefix codes in the continuous form:
@@
with $\lambda>0$. This implies that violating the Kraft–McMillan inequality would imply a penalty on the Lagrangian.
@@
$1=p(\mathcal{C})=\int_{\mathcal{C}}r(x)\nu(dx)=\lambda\int_{\mathcal{C}}\exp(\ell(x))\nu(dx)=\lambda.$ (9)
@@ 2.2 Renormalizing the Codeword Distribution
The process used below is in essence the one of “coarse‐graining” by replacing $x$ with $\log x$ followed by “renormalizing” the resulting probability (or code length). However, it appears to be technically much simpler than the classical renormalization techniques in physics [12, 13, 14] due to its code-theoretical nature.
Thus, put
@@
The above, together with the natural constraint $\ell(x)\geq 0$, implies that $\ell(x)$ is given by the sum of iterated logarithms $\log(\log(\ldots\log(x)\ldots)$ that continues until we obtain a negative number.
@@
We shall call any code satisfying Theorems 2.1 - 2.2 an Elias’ Omega code, for short.
@@ 2.3 Renormalization Flow Universality
where the map $\log:\exp(x)\to x$ implements a coarse‐graining step or scale reduction, while the factor $\exp(-x)$ ensures $\int_{0}^{\infty}\mathcal{R}[p](x)\,dx=1$.
@@
The fixed point $\ell^{*}(x)$ of this flow is exactly the Elias’ code length discussed above:
$\ell_{*}(x)=T[\ell_{*}](x)=\ell_{*}(\log x)+\log x.$ (21)
Recursively unrolling this relation yields exactly
@@
terminating when the next $\log$ argument drops below $1$, i.e. the “0” termination of Elias’ Omega encoding.
@@
where $m=\log^{*}x$, the iterated logarithm of $x$, known for its very low rate of growth. This implies that
$\ell_{*}(x)=O(\log x),$ (24)
all the while being a self-delimiting code unlike the usual binary encoding, which obviously also has $O(\log x)$ code length.
@@
We can summarize our findings as the following propositions.
Proposition 2.3.
Let $\mathcal{C}$ be an Elias’ Omega code with code length $\ell(x)$ for $x\in\mathcal{C}$. Then
$\ell(x)=\log x+\log^{(2)}x+\ldots+\log^{(m)}x,$
with $m=\log^{*}x$, the iterated logarithm of $x$.
Proposition 2.4.
Let $\ell_{0}(x)$ be some non-negative function on the codewords of an Elias’ Omega code $\mathcal{C}$. Then the repeated iterations of the renormalization map $\ell\mapsto\ell\circ\log+\log$ almost converge to the canonical code length. Namely, if $m=\log^{*}x$, let $\ell_{m}=T^{(m)}[\ell]$, and then
$|\ell(x)-\ell_{m}(x)|<O(1)$
independent of $x$.
@@ 2.4 Passing to the Discrete Encoding
Naturally, if we now recast our observation into the discrete setting, Elias’ Omega encoding [1] will provide a practical realization with very close properties.
As it was mentioned before, Elias’ Omega encoding has to be complete. Indeed, let $L(n)$ be the length of $n$ in Elias’ Omega encoding, and let $\beta(n)=\lfloor\log_{2}n\rfloor+1$ be the binary length of $n$. Then we have
$L(n)=L(\beta(n)-1)+\beta(n).$ (28)
@@
$S_{k}=\sum_{n\in I_{k}}2^{-L(n)}=\sum_{n\in I_{k}}2^{-L(k-1)-k}=2^{-L(k-1)-1}.$ (30)
Then for the Kraft’s sum $S=\sum_{n\in\mathbb{N}}2^{-L(n)}$ we get
$S=\sum^{\infty}_{k=1}S_{k}=S_{1}+\sum^{\infty}_{k=2}S_{k}=\frac{1}{2}+\sum^{\infty}_{k=2}2^{-L(k-1)-1}=\frac{1}{2}+\frac{1}{2}\sum^{\infty}_{k=1}2^{-L(k)}=\frac{1+S}{2}.$ (31)
@@ Acknowledgments
3 Appendix: Kraft–McMillan and Quantization
@@ 3.1 Mixed Source Distributions
Instead of the $D$-ary code length $\ell_{D}(x)$ for $x\in\mathcal{C}$, we shall use the baseless (more precisely, base–invariant) code length $\ell(x)=\ell_{D}(x)/\log D$.
@@ 3.3 Prefix-free Quantized Codes
3.4 Heisenberg–style and Boltzmann–style relations
@@
Here $H(\pi)=-\sum_{s}\pi_{s}\ln\pi_{s}$ is the Shannon entropy of the quantized source.
@@ 3.5 Appendix: Shannon’s Entropy Lower Bound
Let $\mathcal{C}\subset[1,+\infty)$ be a code as defined throughout the manuscript, and let its course distribution satisfy

### v4

@@ Abstract
In the present paper we give a derivation of Elias’ Omega code from physics principles by combining a constrained variational formulation of prefix coding with a renormalization flow on codeword distributions.
Starting from a Lagrangian that minimizes average codelength under the Kraft–McMillan constraint, we show that the implied distribution is a fixed point of a coarse–graining map, yielding the canonical iterated log-sum length, asymptotically up to an additive constant.
This establishes completeness and asymptotic optimality, and connects universal integer coding with coarse–grained entropy, uncertainty–type bounds, and entropy relations familiar from statistical physics.
@@ 1 Introduction
Universal integer codes are a classical topic at the interface of information theory and the theory of algorithms. Elias’ family of $\gamma$–, $\delta$–, and $\omega$–codes gives practical schemes whose expected lengths track entropy without prior knowledge of the source law – a property known as universality [1].
The structure of Elias’ $\omega$ exhibits a recursive nature: the length of an integer depends on the size of the integer and on the sizes of the integers describing that size, and so on. It can also be interpreted as scale-invariance, and thus enables a renormalization approach to the probability distribution of codelengths.
Let $s=\text{bin}(x)$ be the binary string representation of a positive decimal number $x$, and let $x=\text{dec}(s)$ be the decimal value corresponding to a binary string $s$, possibly with leading zeros. Then Elias’ $\omega$ is defined by the following two algorithms [2].
Algorithm 1 Elias’ $\omega$: Encoding
@@
Algorithm 2 Elias’ $\omega$: Decoding
@@
The code description given above implies the following curious identity about its codeword lengths.
Lemma 1.1 (Elias $\omega$ codelength).
Let $n_{0}=n$ and recursively define $n_{j+1}=\lfloor\log_{2}n_{j}\rfloor+1$. Stop at the first $m$ such that $n_{m}=1$. Then the Elias $\omega$ codelength is
$\ell_{\omega}(n)\;=\;1+\sum_{j=0}^{m-1}\big(\lfloor\log_{2}n_{j}\rfloor+1\big).$ (1)
Consequently,
$\ell_{\omega}(n)\;=\;\sum_{j=0}^{m-1}\log_{2}n_{j}\;+\;O(\log^{*}n),$ (2)
where $\log^{*}n$ is the iterated base-$2$ logarithm.
The $\omega$–code is known to be prefix-free, meaning that no codeword can represent an initial segment of another codeword. Such codes are uniquely decodable since there is no ambiguity between the source elements encoded in a sequence and the sequence of codewords representing them.
Being prefix-free turns out to be equivalent to the classical Kraft–McMillan [3, 4] conditions on the code word lengths. This condition turns out to be equivalent to a specific bound on the statistical sum with respect to a Gibbs measure.
We recast Elias’ $\omega$–code in a language closer to statistical physics by combining
  1. (i)
    a Lagrangian formulation of optimal prefix coding under the Kraft–McMillan conditions [3, 4, 5], and
  2. (ii)
    a renormalization map implementing coarse-graining of the codeword distribution, and normalization.
The fixed point of the induced flow on codelengths is exactly the $\omega$–codelength; moreover, this fixed point has a wide basin of attraction.
The Lagrangian approach has been used before by Kafri to analyze Benford’s law [6], and by Visser [7] to derive Zipf’s law from the Maximum Entropy Principle [8, 9].
The renormalization approach, to the best of our knowledge, has not appeared in the earlier literature. However, it is very naturally connected to the original definition of Elias’ $\omega$–code and its recursive nature.
The encoding problem becomes a transparent toy model where scale-invariance is enforced by information-theoretical constraints (being uniquely decodable) rather than dynamics, linking two methodological pillars of statistical physics: variational principles and renormalization flows [10, 11, 12].
Moreover, our analysis applies to codes with both discrete and continuous components of their codelength distribution. The usual information constraint for a discrete code that we use is the classical Kraft–McMillan inequality that guarantees the code is uniquely decodable. Reformulating it in terms of a Gibbs measure allows writing down the continuous analogue.
Moreover, a continuous code can be quantized by binning together some of the codewords in order to achieve a series of continuous approximations with varying resolution. At each resolution level, the classical discrete theory applies, while This procedure recovers information-theoretical analogues of both classical Heisenberg’s (resolution–complexity tradeoff in code quantization) and Boltzmann’s (persistence of entropy at each resolution level) relations.
In this work, we achieve several goals:
  • give a variational derivation showing that the optimizer saturates the Kraft–McMillan inequality and identifies the implied codeword probability law – a property otherwise called completeness (Theorem 2.1);
  • quantify asymptotic optimality $\mathbb{E}[\ell]\leq H(p)+O(1)$, with Shannon’s lower bound $\mathbb{E}[\ell]\geq H(p)$, and interpret the discrepancies using physics analogies: code quantization, Heisenberg’s and Boltzmann’s relations (Theorem 2.2);
  • define a simple renormalization operator on codeword probability laws and the induced map on lengths; we prove that its fixed point is the iterated-log form of $\omega$ and that the flow is attractive up to a reasonably small error (Theorem 2.3).
@@ 2.1 Lagrangian Optimization for Codelength
i.e. a weighted sum of delta-functions together with a continuous probability density. Here $W$ is at most countable, $\rho\geq 0$ is integrable, $W\cap\mathrm{supp}\,\rho=\emptyset$, and
@@
Let us recast the Kraft–McMillan inequality [3, 4] for prefix codes in the continuous form:
@@
with $\lambda>0$. This implies that violating the Kraft–McMillan inequality would impose a penalty on the Lagrangian.
@@
$1=p(\mathcal{C})=\int_{\mathcal{C}}r(x)\nu(dx)=\lambda\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)=\lambda.$ (11)
@@ 2.2 Renormalizing the Codeword Distribution
The process used below is in essence the one of coarse‐graining by replacing $x$ with $\log x$ followed by renormalizing the resulting probability (or codelength). However, it appears to be technically much simpler than the classical renormalization techniques in physics [10, 11, 12] due to its code-theoretical nature.
Let us put
@@
The above, together with the natural constraint $\ell(x)\geq 0$, implies that $\ell(x)$ is given by the sum of nested logarithms $\log(\log(\ldots\log(x)\ldots)$ that continues until we obtain a negative number.
@@
We shall call any code satisfying Theorems 2.1 - 2.2 an $\omega$–type code, for short.
@@ 2.3 Renormalization Flow Universality
where the map $\log:\exp(x)\to x$ implements a coarse‐graining step or scale reduction, while the factor $\exp(-x)$ ensures
$\int_{1}^{\infty}\mathcal{R}[p](x)\,dx=\int^{\infty}_{0}\mathcal{R}[p](\exp(x))=$
$=\int^{\infty}_{0}p(x)\exp(-x)\cdot\exp(x)dx=$
$=\int^{1}_{0}0\cdot dx+\int^{\infty}_{1}p(x)dx=1,$
where we assume $p(x)\equiv 0$ outside of the usual codeword range $1,+\infty)$.
@@
The fixed point $\ell^{*}(x)$ of this flow is almost the Elias’ codelength from Lemma 1.1. Namely, we have
$\ell_{*}(x)=\mathcal{T}[\ell_{*}](x)=\ell_{*}(\log x)+\log x,$ (23)
and recursively unrolling this relation yields
@@
terminating when the next $\log$ argument drops below $1$, which is directly analogous to the “0” termination of Elias’ $\omega$ encoding.
If we rewrite Elias’ $\omega$ codelength using natural logarithms instead of base-$2$ logarithms111i.e. we measure codelengths in nats rather than in bits., we get
$\ell_{\omega}(x)=\ell_{*}(x)+O(\log^{*}x)=\ell_{*}(x)+o(\log(x)^{\varepsilon}),$ (25)
where $\log^{*}(x)$ is the iterated logarithm of $x$ known for its very slow growth that guarantees the above equality for any $\varepsilon>0$.
@@
with $m=\log^{*}x$. This, together with the obvious $\ell_{*}(x)\geq\log x$, implies that
$\ell_{*}(x)=\Theta(\log x),$ (27)
all the while being a self-delimiting code unlike the usual binary encoding, which obviously also has $\Theta(\log x)$ codelength.
@@
We can summarize our findings as the following theorem.
Theorem 2.3.
Let $\ell_{0}(x)$ be some non-negative function on the codewords of an $\omega$–type code $\mathcal{C}$. Then the repeated iterations of the renormalization map $\ell\mapsto\ell\circ\log+\log$ almost converge to the canonical codelength. Namely, let $m=\log^{*}x$ and $\ell_{m}=\mathcal{T}^{(m)}[\ell]$. Then
$|\ell_{w}(x)-\ell_{m}(x)|=\mathrm{const}+o(\ell_{\omega}(x)^{\varepsilon}),$
for any $\varepsilon>0$.
@@ 2.4 Passing to the Discrete Encoding
Naturally, if we now recast our observation into the discrete setting, Elias’ $\omega$ encoding [2] will provide a practical realization with very close properties.
As it was mentioned before, Elias’ $\omega$ encoding has to be complete. Indeed, let $\ell_{\omega}(n)$ be the length of $n$ in Elias’ $\omega$ code, and let $\beta(n)=\lfloor\log_{2}n\rfloor+1$ be the binary length of $n$. Then we have
$\ell_{\omega}(n)=\ell_{\omega}(\beta(n)-1)+\beta(n).$ (31)
@@
$S_{k}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(n)}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(k-1)-k}=2^{-\ell_{\omega}(k-1)-1}.$ (33)
Then for the Kraft’s sum $S=\sum_{n\in\mathbb{N}}2^{-\ell_{\omega}(n)}$ we get
$S=\sum^{\infty}_{k=1}S_{k}=S_{1}+\sum^{\infty}_{k=2}S_{k}=\frac{1}{2}+\sum^{\infty}_{k=2}2^{-\ell_{\omega}(k-1)-1}=\frac{1}{2}+\frac{1}{2}\sum^{\infty}_{k=1}2^{-\ell_{\omega}(k)}=\frac{1+S}{2}.$ (34)
@@ Acknowledgments
3 Appendix: Continuous codes and Quantization
@@ 3.1 Mixed Source Distributions
Instead of the $D$-ary codelength $\ell_{D}(x)$ for $x\in\mathcal{C}$, we shall use the baseless (more precisely, base-invariant) codelength $\ell(x)=\ell_{D}(x)/\log D$.
@@ 3.3 Prefix-free Quantized Codes
3.4 Heisenberg-style and Boltzmann-style relations
@@
Here
$H(\pi)=-\sum_{s}\pi_{s}\ln\pi_{s}$
is the Shannon entropy of the quantized source, and
$D_{\mathrm{KL}}(\pi\|q)=\sum_{s}\pi_{s}\log\frac{\pi_{s}}{q_{s}}$
is the Kullback–Leibler (KL) divergence of two distributions [5].
@@ 3.5 Appendix: Shannon’s Entropy Lower Bound
Let $\mathcal{C}\subset[1,+\infty)$ be a code as defined throughout the manuscript, and let its coarse distribution satisfy